What is the arithmetic mean of first 16 natural number with weights being the number itself?

Grade:12th pass

1 Answers

Vidhi Shah
6 years ago
The sum of first n natural numbers is given by the formula, Sn=n(n+1)2Put n=16 to get the sum of first 16 natural numbers. S16=16 (16+1)2 =16×172 =8×17 =136The required arithmetic mean is given by, mean=sum of obervationsnumber of observations =13616 =8.5